OpenStudy (anonymous):

Which is the simplified form of the product of 18 a squared b to the fifth all over the product of 6 a to the eighth b

OpenStudy (anonymous):

Hint: \(\Large \dfrac{18a^2b^5}{6a^8b}=\dfrac{18}{6}a^{2-8}b^{5-1}\)